# Transmission lines; Defect detection; UAV imagery; Power industry; Transformer-based detectors

### v2

@@ 3.5 Focaler-Wise-SIoU Loss Function
In the task of transmission line defect detection, where the majority of objects are small and often exhibit subtle spatial differences, accurate bounding box regression plays a crucial role. IoU-based loss functions have become the de facto standard for this purpose in modern object detectors, yet they still suffer from several inherent limitations. The vanilla IoU loss only considers the overlap ratio and fails to capture spatial misalignment when boxes do not intersect. Its variants, such as GIoU(Rezatofighi et al., 2019), DIoU(Zheng et al., 2019), and CIoU(Zheng et al., 2021), partially alleviate this issue by introducing penalties related to enclosure area, center distance, or aspect ratio. However, these approaches remain limited in their ability to jointly model geometric alignment and shape consistency. More advanced designs such as SIoU(Gevorgyan, 2022) attempt to incorporate angular, distance, and shape discrepancies, providing richer geometric information. Nevertheless, they still treat all samples equally, regardless of their difficulty, which results in an imbalanced gradient contribution. Easy samples dominate the optimization, while hard samples, particularly those corresponding to small defects with low signal-to-noise ratios, receive insufficient emphasis. Furthermore, in high-IoU regimes where precise localization is critical for tiny defect regions, conventional losses often lack discriminative power, leading to slower convergence and degraded accuracy.
@@ 3.6 CSG-ADCD
NO Label Name NO Label Name
0 zcjyz Normal Insulator 1 jyzwh Polluted Insulator
2 dgss Damaged Pole 3 zxqs Missing Tie Wire
4 zxst Loose Tie Wire 5 jyzsl Insulator Flashover
6 nw Bird Nest 7 jyzps Broken Insulator
@@ 4.1 Model Training and Evaluation Metrics
4.2 Comparison of attention maps
To intuitively demonstrate the effectiveness of the proposed modules, we generate heatmaps from the baseline RT-DETR model and the model with the proposed modules (EEBlock, SPD Convolution, and CSDMAM) on the same input image. These heatmaps illustrate the attention distribution of the models during the recognition process. As shown in Fig.6, subfigures (a), (c), and (e) represent the attention maps from the baseline model corresponding to the modules before integration, while subfigures (b), (d), and (f) show the attention maps after incorporating the proposed modules. It can be observed that the attention distribution after the inclusion of the modules is more fine-grained, and the target contours are more distinct, indicating that the proposed modules enhance the model’s ability to preserve more detailed information.
@@ 4.3 Ablation Study of the Proposed Method
To analyze the impact of each innovative component in CSG-ADCD, we conducted ablation experiments by systematically including and excluding four key modules: EER, SPD, CSDMAM, and FWS Loss. The results, shown in the Table 2, demonstrate the contribution of each module to the model’s overall performance across various evaluation metrics.
@@
EER SPD CSDMAM FWS Loss GFLOPs Precision Recall $\mathrm{mAP}_{50}$ $\mathrm{AP}_{s}$
$\checkmark$ $\times$ $\times$ $\times$ $57.0$ $0.369$ $0.177$ $0.163$ $0.071$
@@ 4.4 Comparisons With Previous Methods
Before delving into aggregate statistics, we first compare the recognition performance of TinyDef-DETR with that of other representative models through illustrative case studies.
Figure 7: (a)–(e)
Table LABEL:tab:model_performance summarizes the overall performance of representative detectors on the CSG-ADCD dataset. The comparison covers lightweight and standard variants of YOLO (v5, v8, v10, v11), Transformer-based models such as RT-DETR and Conditional-DETR, as well as classical baselines including RetinaNet and DETR. Key indicators, including the number of parameters, computational complexity (GFLOPs), Precision, Recall, $\text{mAP}{50}$, and $\text{mAP}{50:95}$, are reported to provide a comprehensive view of accuracy–efficiency trade-offs.
@@
Table 3: Model Performance Evaluation: Precision, Recall, and mAP for CSG-ADCD
@@
Conditional-DETR-R18(Meng et al., 2023) 0.150 0.073 0.082 0.009 0.016
@@ 4.5 Generalization Experiments for the Proposed Methods
Since small objects account for a large proportion in this dataset, to further validate the generalization capability of the proposed method, we additionally conduct generalization experiments on the VisDrone(Zhu et al., 2022) dataset. VisDrone is a large-scale benchmark dataset specifically constructed for visual perception tasks under the perspective of unmanned aerial vehicles (UAVs). Created by the Aerial Vision Group at the University of Science and Technology of China, the dataset comprises over 26,000 high-resolution images and 263 video sequences, captured across diverse urban and suburban environments such as streets, rooftops, playgrounds, and traffic scenes. It includes more than 100,000 manually annotated object instances spanning 10 common categories, including pedestrians, cars, buses, trucks, bicycles, and others. Each object is labeled with rich metadata, including bounding boxes, object class, occlusion level, and truncation ratio. Due to its dense object distribution, high proportion of small objects, severe occlusion, and large-scale variation in viewpoint and illumination, VisDrone presents considerable challenges for detection and tracking models. Consequently, it has been widely adopted as a standard benchmark to evaluate the robustness and generalization ability of UAV-based visual recognition algorithms in real-world complex scenarios. To further assess the generalization capability of the proposed TinyDef-DETR, we conducted experiments on the VisDrone dataset and compared its performance against representative state-of-the-art YOLO series models, FFCA-YOLO, and its baseline RT-DETR-R18. The results are summarized in Table 4.5.
@@
Cascade-RCNN-R50(Cai and Vasconcelos, 2017) $69.3$ $236.0$ $0.326$ $0.197$ $0.099$ $0.309$
@@
5 Discussion
This study set out to improve small-defect detection for UAV-based transmission line inspection by addressing three well-known bottlenecks of mainstream detectors: the loss of discriminative details caused by strided downsampling, the insufficient boundary sensitivity of lightweight CNN backbones, and the limited ability of current architectures to jointly model global context with fine-grained local cues in cluttered backgrounds. Building on the RE-DETR framework, the proposed TinyDef-DETR introduces EEConv-based detail enhancement, a stride-free space-to-depth (SPD) downsampling strategy, a cross-stage dual-domain multi-scale attention module (CSDMAM), and a Focaler-Wise-SIoU (FWS) regression loss tailored to small objects. Across the imbalanced and small-object-dominated CSG-ADCD dataset, TinyDef-DETR achieves consistent improvements over competitive baselines while incurring only modest computational overhead, and its effectiveness further generalizes to the VisDrone benchmark.
The experimental results confirm several of our initial hypotheses. First, preserving spatial detail is of paramount importance for small-defect detection. The SPD module, which rearranges spatial information into channel dimensions before any stride is applied, significantly improves recall and small-object average precision by preventing the early loss of discriminative micro-structures. Second, embedding gradient priors into the backbone through EEConv enhances boundary localization. By reintroducing directional gradient sensitivity in a re-parameterizable form, EEConv captures sharper contours of insulators and fittings, leading to higher precision and more stable training without additional inference cost. Third, the proposed CSDMAM demonstrates that global–local dual-domain attention is especially beneficial when upstream features retain pixel-level details and edge cues. Its combination of large-kernel depthwise convolutions for elongated structures and frequency-domain reweighting for texture calibration effectively complements the detail-preserving modules. Finally, the introduction of the FWS regression loss proves essential for handling small objects, as it strengthens optimization in the high-IoU regime and reallocates learning capacity toward moderately hard but informative samples, yielding notable gains in localization accuracy.
Quantitatively, TinyDef-DETR improves $\mathrm{mAP}_{50}$ on CSG-ADCD from 0.163 to 0.275, representing a relative gain of 68.7%, while recall increases from 0.177 to 0.263. The improvement in $\mathrm{AP}_{s}$ from 0.074 to 0.119 further underscores the model’s advantage on small defects. These gains come with a moderate increase in computation from 57.0 to 65.3 GFLOPs, which is acceptable in practical scenarios. Generalization experiments on VisDrone also validate the proposed design, with $\mathrm{mAP}_{50}$ rising from 0.333 to 0.372 and $\mathrm{AP}_{s}$ from 0.139 to 0.148, suggesting that the enhancements address dataset-agnostic weaknesses in small-object detection rather than overfitting to a specific domain.
Beyond quantitative gains, the study highlights several practical implications for UAV-based inspection. The re-parameterized EEConv provides edge-aware yet deployment-friendly backbones, ensuring inference simplicity for embedded UAV processors. The SPD strategy demonstrates that lossless early downsampling should become a standard design choice for aerial small-object detection, as it improves recall with negligible computational burden. Moreover, CSDMAM proves particularly effective for line-structured scenes, as its anisotropic receptive fields and frequency calibration align well with the geometric characteristics of power lines and insulators.
Nevertheless, the approach has several limitations. The annotations in CSG-ADCD are subject to class imbalance and potential bias, as some rare defect types remain underrepresented. Furthermore, the evaluation focuses on UAV imagery with relatively consistent flight altitudes and backgrounds, leaving the robustness of the method under different sensing modalities, weather conditions, or severe motion blur less certain. Although $\mathrm{AP}_{s}$ improves substantially, the absolute performance on ultra-small or heavily occluded instances remains modest, highlighting the difficulty of this task. In addition, the computational cost, while moderate on GPUs, may still present challenges for real-time onboard deployment on low-power UAV hardware.
Future work should therefore consider extending the framework along several directions. Incorporating temporal and multi-view reasoning may exploit redundancy across inspection passes to disambiguate subtle defects. Pretraining with self-supervised or foundation-model approaches on large-scale aerial corpora could further improve generalization. Structural priors derived from the physics and geometry of transmission components offer a promising avenue to reduce false positives. Coupling detection with segmentation or keypoint localization could sharpen predictions for minute parts, while semi-supervised and active learning strategies would help leverage abundant unlabeled inspection video. Finally, model compression techniques such as pruning, quantization, or hardware-aware architecture search will be necessary to enable real-time UAV deployment without sacrificing accuracy.
In conclusion, the results substantiate our central hypothesis: when early spatial detail is preserved through SPD, edge sensitivity is enhanced by EEConv, global–local dual-domain context is introduced with CSDMAM, and regression is guided by a difficulty-aware FWS loss, DETR-style detectors achieve markedly stronger performance on small, ambiguous defects in aerial inspection. The proposed design is modular, training-efficient, and deployment-conscious, offering a practical pathway toward reliable and scalable UAV-based transmission line monitoring.
@@ 6 Conclusions
This work proposed TinyDef-DETR, a detection framework tailored for small-defect inspection in UAV-based power line imagery. By integrating a stride-free space-to-depth downsampling strategy, an edge-enhanced convolution for boundary-aware representation, a cross-stage dual-domain multi-scale attention module, and a difficulty-adaptive regression loss, the method addresses key limitations of mainstream detectors. Extensive experiments on the CSG-ADCD dataset demonstrate substantial improvements in both precision and recall, particularly for small objects, while maintaining moderate computational requirements. Further validation on the VisDrone benchmark confirms that the proposed components generalize beyond the specific domain of transmission line inspection.
The study highlights that preserving fine spatial details, embedding gradient priors, and combining global context with local edge cues are critical for reliable detection of small and ambiguous targets. Although challenges remain—such as class imbalance, domain generalization, and onboard deployment constraints—the presented framework offers a practical and extensible foundation for UAV-assisted power grid monitoring. Future directions include leveraging temporal and multi-view information, incorporating self-supervised pretraining, exploring structural priors, and advancing model compression techniques to achieve real-time performance on resource-constrained platforms.

### v3

@@ 3.5 Focaler-Wise-SIoU Loss Function
In the task of transmission line defect detection, where the majority of objects are small and often exhibit subtle spatial differences, accurate bounding box regression plays a crucial role. IoU-based loss functions have become the de facto standard for this purpose in modern object detectors, yet they still suffer from several inherent limitations. The vanilla IoU loss only considers the overlap ratio and fails to capture spatial misalignment when boxes do not intersect. Its variants, such as GIoU(Rezatofighi et al., 2019), DIoU(Zheng et al., 2020), and CIoU(Zheng et al., 2022), partially alleviate this issue by introducing penalties related to enclosure area, center distance, or aspect ratio. However, these approaches remain limited in their ability to jointly model geometric alignment and shape consistency. More advanced designs such as SIoU(Gevorgyan, 2022) attempt to incorporate angular, distance, and shape discrepancies, providing richer geometric information. Nevertheless, they still treat all samples equally, regardless of their difficulty, which results in an imbalanced gradient contribution. Easy samples dominate the optimization, while hard samples, particularly those corresponding to small defects with low signal-to-noise ratios, receive insufficient emphasis. Furthermore, in high-IoU regimes where precise localization is critical for tiny defect regions, conventional losses often lack discriminative power, leading to slower convergence and degraded accuracy.
@@ 4.1 CSG-ADCD
NO Label Name
0 zcjyz Normal Insulator
1 jyzwh Polluted Insulator
2 dgss Damaged Pole
3 zxqs Missing Tie Wire
4 zxst Loose Tie Wire
5 jyzsl Insulator Flashover
6 nw Bird Nest
7 jyzps Broken Insulator
@@ 4.2 Model Training and Evaluation Metrics
4.3 Comparison of heatmaps
To intuitively demonstrate the effectiveness of the proposed modules, we generate heatmaps from the baseline RT-DETR model and the model with the proposed modules (EE-ResNet, SPD Convolution, and CSDMAM) on the same input image. These heatmaps illustrate the attention distribution of the models during the recognition process. As shown in Fig.7, subfigures (a), (c), and (e) represent the attention maps from the baseline model corresponding to the modules before integration, while subfigures (b), (d), and (f) show the attention maps after incorporating the proposed modules. It can be observed that the attention distribution after the inclusion of the modules is more fine-grained, and the target contours are more distinct, indicating that the proposed modules enhance the model’s ability to preserve more detailed information.
@@ 4.4 Ablation Study of the Proposed Method
To analyze the impact of each innovative component in CSG-ADCD, we conducted ablation experiments by systematically including and excluding four key modules: EER (Edge-Enhanced ResNet), SPD (Stride-free Space-to-Depth), CSDMAM (Cross-Stage Dual-Domain Multi-Scale Attention Module), and FWS Loss (Focaler-Wise-SIoU Loss). The results, shown in Table 2, demonstrate the contribution of each module to the model’s overall performance across various evaluation metrics.
@@
EEREER SPDSPD CSDMAMCSDMAM FWSLossFWSLoss GFLOPs Precision Recall $\mathrm{mAP}_{50}$ $\mathrm{AP}_{s}$
$\times$ $\times$ $\times$ $\times$ $57.0$ $0.369$ $0.177$ $0.163$ $0.071$
@@ 4.5 Comparisons With Previous Methods
Before presenting the experimental data, to visually highlight the differences between TinyDef-DETR and other representative detectors, we conduct a case study on four representative images from the CSG-ADCD dataset. To save space and facilitate readability in a scenario where the majority of targets are small objects, we render the detection bounding boxes with filled colors rather than outlines only. This visualization provides a clearer comparison of recognition results across different models, and the detection results are illustrated in Fig. 8.
Figure 8: Comparison of detection results from four object detection methods on four representative test samples.Rows (a)–(d) correspond to different test images, and the four columns show the results obtained by TinyDef-DETR, DINO, RT-DETR-R18, and YOLO v5s, respectively.
Table LABEL:tab:model_performance summarizes the overall performance of representative detectors on the CSG-ADCD dataset. The comparison covers lightweight and standard variants of YOLO (v5, v8, v10, v11), Transformer-based models such as RT-DETR and Conditional-DETR, as well as classical baselines including RetinaNet and DETR. Key indicators, including the number of parameters, computational complexity (GFLOPs), Precision, Recall, $\text{mAP}_{50}$, and $\text{mAP}_{50:95}$, are reported to provide a comprehensive view of accuracy–efficiency trade-offs. The best values in each column are highlighted in bold for clarity.
@@
It is also worth noting that although DINO achieves slightly higher performance than TinyDef-DETR on certain metrics, its training requires approximately 79.8 hours, which is significantly longer. Such excessive training time limits its practical applicability in real-world UAV-based inspection tasks, whereas TinyDef-DETR achieves a more balanced trade-off between accuracy and efficiency.
Table 3: Model Performance Evaluation: Precision, Recall, and mAP for CSG-ADCD (best results highlighted in bold).
@@
Conditional-DETR-R18(Meng et al., 2021) 0.150 0.073 0.082 0.009 0.016
@@ 4.6 Generalization Experiments for the Proposed Methods
Since small objects account for a large proportion in this dataset, to further validate the generalization capability of the proposed method, we additionally conduct generalization experiments on the VisDrone(Zhu et al., 2022) dataset. VisDrone is a large-scale benchmark dataset specifically constructed for visual perception tasks under the perspective of unmanned aerial vehicles (UAVs). Created by the Aerial Vision Group at the University of Science and Technology of China, the dataset comprises over 26,000 high-resolution images and 263 video sequences, captured across diverse urban and suburban environments such as streets, rooftops, playgrounds, and traffic scenes. It includes more than 100,000 manually annotated object instances spanning 10 common categories, including pedestrians, cars, buses, trucks, bicycles, and others. Each object is labeled with rich metadata, including bounding boxes, object class, occlusion level, and truncation ratio. Due to its dense object distribution, high proportion of small objects, severe occlusion, and large-scale variation in viewpoint and illumination, VisDrone presents considerable challenges for detection and tracking models. Consequently, it has been widely adopted as a standard benchmark to evaluate the robustness and generalization ability of UAV-based visual recognition algorithms in real-world complex scenarios. To further assess the generalization capability of the proposed TinyDef-DETR, we conducted experiments on the VisDrone dataset and compared its performance against representative state-of-the-art YOLO series models, FFCA-YOLO, and its baseline RT-DETR-R18. The results are summarized in Table 4.6, where the best values in each column have been highlighted in bold for clarity.
@@
Cascade-RCNN-R50(Cai and Vasconcelos, 2021) $69.3$ $236.0$ $0.326$ $0.197$ $0.099$ $0.309$
@@ 5 Conclusions
In this paper, we proposed TinyDef-DETR, a DETR-based framework tailored for defect detection in UAV imagery of transmission lines. By integrating an edge-enhanced ResNet backbone, a stride-free space-to-depth module, a cross-stage dual-domain multi-scale attention mechanism, and a Focaler-Wise-SIoU regression loss, the model effectively addresses the challenges posed by small, ambiguous, and background-cluttered defects.
Extensive experiments on the CSG-ADCD dataset, as well as real-world UAV imagery, demonstrated that TinyDef-DETR consistently outperforms representative detectors in terms of detection accuracy, recall, and robustness, while maintaining computational efficiency. Ablation studies further verified the effectiveness of each proposed component, highlighting their complementary contributions to the overall performance.
The results confirm that TinyDef-DETR not only achieves strong detection capability for small and difficult targets but also provides a practical and efficient solution for real-world UAV-based inspection scenarios. Given its modular and lightweight design, the proposed framework holds promise for broader applications in power system inspection and can be extended to other small-object detection tasks in aerial remote sensing.